### v1

# Climbing the Ladder of Interpretability with Counterfactual Concept Bottleneck Models

Abstract
Current deep learning models are not designed to simultaneously address three fundamental questions: predict class labels to solve a given classification task (the “What?”), explain task predictions (the “Why?”), and imagine alternative scenarios that could result in different predictions (the “What if?”). The inability to answer these questions represents a crucial gap in deploying reliable AI agents, calibrating human trust, and deepening human-machine interaction. To bridge this gap, we introduce CounterFactual Concept Bottleneck Models (CF-CBMs), a class of models designed to efficiently address the above queries all at once without the need to run post-hoc searches. Our results show that CF-CBMs produce: accurate predictions (the “What?”), simple explanations for task predictions (the “Why?”), and interpretable counterfactuals (the “What if?”). CF-CBMs can also sample or estimate the most probable counterfactual to: (i) explain the effect of concept interventions on tasks, (ii) show users how to get a desired class label, and (iii) propose concept interventions via “task-driven” interventions.
1 Introduction
To calibrate human trust and enhance human-machine interactions, deep learning (DL) models should learn how to master three fundamental questions: predict class labels for new inputs (the “What?”), explain predictions (the “Why?”), and imagine alternative scenarios that could result in different predictions (the “What if?”)—also known as counterfactual explanations (Wachter et al., 2017). We refer to this escalating sequence of questions as the “ladder of interpretability” (Figure 1), taking inspiration from a similar analogy proposed by Pearl et al. (2016). Crucially, no DL model is currently designed to cover the whole ladder.
Figure 1: CF-CBMs are designed to address three fundamental questions represented by the “ladder of interpretability”: 1. predict class labels for a given task (“What?”), 2. explain predictions via learned concepts (“Why?”), and 3. imagine concept-based counterfactuals that would have led to different class predictions (“What if?”).
At the base of the ladder, standard DL models excel at addressing “what?” questions. Standing on the second rung, Concept Bottleneck Models (CBMs) (Koh et al., 2020) currently represent a higher-ranked DL model in this ladder, as they are designed to answer both “what?” and “why?” questions. To this end, CBMs learn to explain their own predictions by first predicting a set of human-understandable concepts (Kim et al., 2018) and then using the learned concepts to solve the task. Yet, their design and training do not extend to “what if?” queries, leaving the ladder’s third rung still unreachable for these models.
Contributions.
To bridge this gap, we introduce CounterFactual Concept Bottleneck Models (CF-CBM), the first CBMs designed to cover the entire ladder of interpretability. Our key innovation lies in a latent process generating two similar concept vectors. The model uses the first concept vector to predict the target class label (similar to standard CBMs), while the second one is used to predict a different class label. Our experimental results show that CF-CBMs generate: (“What?”) accurate concept/class predictions, (“Why?”) simpler and more robust explanations compared to standard CBMs, (“What if?”) interpretable concept-based counterfactuals. Compared to standard CBMs, CF-CBMs can also sample or estimate the most probable counterfactual to: (i) explain the effect of concept interventions on the classification task, (ii) show users how to get a desired class label, and (iii) propose concept interventions via “task-driven” interventions. The code is available at https://github.com/gabriele-dominici/Counterfactual-CBM.
Figure 2: Counterfactual CBMs inference. For a given input sample, the task predictor answers “what?” queries predicting class labels. The concept predictor answers “why?” queries explaining class predictions via concept labels when taking as input the latent concept distribution $z$. When taking the counterfactual latent distribution $z^{\prime}$, the concept predictor answers “what if?” queries via concept-based counterfactuals.
2 Background
Concept Bottleneck Models (CBMs).
A CBM (Koh et al., 2020; Poeta et al., 2023) is a function composed by: (i) a concept encoder $g:X\to C$ mapping each entity $x\in X\subseteq\mathbb{R}^{d}$ (e.g., an image) to a set of $k$ concepts $\hat{c}\in C\subseteq[0,1]^{k}$ (e.g., “red”, “round”), and (ii) an interpretable (Rudin, 2019) task predictor $f:C\to Y$ mapping concepts to a set of $m$ classes $\hat{y}\in Y\subseteq[0,1]^{m}$ (e.g., “apple”) representing a downstream task. Each component $g_{i}$ and $f_{j}$ denotes the prediction score of the $i$-th concept and $j$-th class, respectively. During training, concept and class predictions $(\hat{c},\hat{y})$ are aligned with ground-truth labels $(c,y)$. At test time, a CBM provides explanations for class predictions indicating the presence or absence of a concept, and human experts may intervene on mispredicted concept labels to improve CBMs’ task performance. In a probabilistic perspective (Bahadori and Heckerman, 2020; Misino et al., 2022), (ideal) concepts $c$ represent key factors of variation (Kingma and Welling, 2013; Zarlenga et al., 2023) of observed data $x$ and $y$, as illustrated in the graphical model $x$$c$$y$
Counterfactual Explanations.
A counterfactual is a hypothetical statement in contrast with actual events that helps us understand the potential consequences of different choices or circumstances  (Pearl et al., 2016). In the context of machine learning, Wachter et al. (2017) define counterfactual explanations as an optimization problem representing an answer to a “what if” question. The optimization aims to find for each observation $x$ the closest datapoint $x^{\prime}$ such that a classifier $m:X\rightarrow Y$ produces a class label $m(x^{\prime})$ that is different from the original label $m(x)$.
3 Counterfactual Concept Bottleneck Models
To cover the ladder of interpretability, we propose CounterFactual Concept Bottleneck Models (CF-CBMs, Figure 2), a novel class of interpretable models designed to generate counterfactuals via variational inference. In the following, we motivate CF-CBMs’ design (Sec. 3.1), architecture and optimization objective (Sec. 3.2). Sections 3.4 and 3.3 show how CF-CBMs climb the ladder of interpretability and how users can act upon generated counterfactuals.
3.1 Motivation
We aim to design a model that generates interpretable “concept-based” counterfactuals i.e., counterfactuals enabling humans to act upon meaningful and interpretable units of information (Abid et al., 2022). For instance, “change the intensity of pixel 42 from 0 to 1” does not provide a clear (actionable) explanation of how to transform a “banana” to an “apple” for a human. In contrast, “change the shape of the object to round” provides an interpretable abstraction in terms of human-understandable concepts (Kim et al., 2018). In general, whenever the input features are too fine-grained (e.g., pixel intensities), the corresponding counterfactuals might not follow human semantics (Kim et al., 2018). Concept-based counterfactuals, instead, provide meaningful insights for users to act upon. This makes CF-CBM an ideal architecture to cover the three rungs of the ladder of interpretability.
3.2 Generating Concept-Based Counterfactuals
To realize concept-based counterfactuals, we extend the graphical model $x$$c$$y$ introducing two additional variables: $c^{\prime}$, representing concept-based counterfactual labels; and $y^{\prime}$, representing the counterfactual class label. We represent the counterfactual dependency on the actual concepts $c$ with an arrow from $c$ to $c^{\prime}$, while $y^{\prime}$ depends on $c^{\prime}$:
$x$$c$$y$$c^{\prime}$$y^{\prime}$
We also notice that multiple counterfactuals can be used to explain any given fact (Wachter et al., 2017; Pearl, 2000). To address this, an ideal solution would be to use a generative approach modeling the discrete distributions over $c$ and $c^{\prime}$. However, modeling such distributions directly might be unfeasible in practice (Richardson and Domingos, 2006) as it requires either modelling complex concept dependencies with discrete distributions which scale exponentially with the number of concepts, or assuming that concepts are independent of each other, an often unrealistic assumption. For this reason, we use a latent variable approach (Kingma and Welling, 2013) to model concept dependencies in a continuous latent distribution, which allows us to model discrete concept values as independent of each other given the latent variable.
Architecture.
CF-CBMs are latent variable models generating counterfactuals via variational inference. To this end, they add two random variables $z$ and $z^{\prime}$ to Diagram 1. These variables represent latent factors of variation whose probability distributions are easier to model and sample compared to those for $c$ and $c^{\prime}$. We also include arrows from $c$ and $y$ to the counterfactual latent distribution $z^{\prime}$ in order to explicitly model the dependency of $z^{\prime}$ on symbolic values of actual concept and class labels, resulting in the following overall probabilistic graphical model:
$x$$c$$y$$c^{\prime}$$y^{\prime}$$z$$z^{\prime}$
This way, the generative distribution factorizes as:
$p(c,y|z)=p(y|c)p(c|z)$ (4)
$p(c^{\prime},y^{\prime}|z^{\prime})=p(y^{\prime}|c^{\prime})p(c^{\prime}|z^{\prime})$ (5)
$p(\mathbf{z}|c,y)=p(z)p(z^{\prime}|z,c,y)$ (6)
In our approach, $p(y|c)$ and $p(y^{\prime}|c^{\prime})$ are modeled as categorical distributions parametrized by the same task predictor $f$; $p(c|z)$ and $p(c^{\prime}|z^{\prime})$ as sets of Bernoulli distributions parametrized by the same concept predictor $\phi_{z}$. In practice, we assume that the input $x$ is always observed at test time, making the term and $p(x|z)$ irrelevant. Finally, $p(z)$ is a standard normal prior distribution and $p(z^{\prime}|z,c,y)$ is a learnable normal prior whose mean and variance are parametrized by a pair of neural networks $\phi_{p\mu}$ and $\phi_{p\sigma}$.
Amortized inference.
CF-CBMs amortize inference needed for training by introducing two approximate Gaussian posteriors $q(z|x)$ and $q(z^{\prime}|z,c,y,y^{\prime})$ whose mean and variance are parametrized by a pair of neural networks $(\phi_{\mu},\phi_{\sigma})$ ($(\phi_{\mu^{\prime}},\phi_{\sigma^{\prime}})$, respectively). The corresponding inference graphical model (i.e. the encoder) is shown in Appendix A.
Optimization problem.
CF-CBMs are trained to maximize the log-likelihood of tuples $(c,y,y^{\prime})$, while observing $x$. Following a variational inference approach, we optimize the evidence lower bound of the log-likelihood, which results (see Appendix A) in the following objective function to maximize:
$\mathcal{L}=$ $\overbrace{\mathbb{E}_{z\sim q(z|x)}[\log p(c|z)]+\log p(y|c)}^{\text{ reconstruction of $c$ and $y$}}-\overbrace{D_{KL}[q(z|x)||p(z)]}^{\text{ prior regularization on $z$}}$
$+$ $\overbrace{\mathbb{E}_{z,z^{\prime},c^{\prime}\sim p(c^{\prime}|z^{\prime})q(z^{\prime}|\alpha)q(z|x))}[\log p(y^{\prime}|c^{\prime})]}^{\text{reconstruction of $y^{\prime}$}}$
$-\overbrace{D_{KL}[q(z^{\prime}|\alpha)||p(z^{\prime}|z,c,y)]}^{\text{prior regularization on $z^{\prime}$ }}$
where $D_{KL}$ is the Kullback–Leibler divergence and $\alpha=(z,c,y,y^{\prime})$. Moreover, in order to enforce concept-based counterfactuals to be as close as possible to the current concept labels, we add an additional term to the objective:
3.3 Climbing the “Ladder of Interpretability”
We claim that CF-CBMs climb the ladder of interpretability by design, answering “what?”, “why?” and “what if?” queries through the following steps (Figure 2):
  1. 1.
    What? predict concept and class labels:
    1. (a)
      Sample from latent concept posterior:
      $z\sim\mathcal{N}(\phi_{\mu}(x),\phi_{\sigma}(x))$
    2. (b)
      Sample to predict concept and class labels:
      $\hat{y}\sim\text{Categorical}(f(\hat{c})),\ \hat{c}\sim\text{Bernoulli}(\phi_{z}(z))$
  2. 2.
    Why? explain class predictions using the interpretable model $f$ which identifies relevant concepts.
  3. 3.
    What if? imagine interpretable counterfactuals via the conditional generation of new concept tuples:
    1. (a)
      Sample from latent counterfactual posterior:
      $z^{\prime}\sim\mathcal{N}(\phi_{\mu^{\prime}}(\alpha),\phi_{\sigma^{\prime}}(\alpha))$
      $\alpha=(z,\hat{c},\hat{y},y^{\prime})$
      $y^{\prime}\sim\text{Categorical}(v),\ v\sim\mathcal{U}\{0,|y|\}$
    2. (b)
      Sample to predict counterfactual labels:
      $\hat{y}^{\prime}\sim\text{Categorical}(f(\hat{c}^{\prime})),\ \hat{c}^{\prime}\sim\text{Bernoulli}(\phi_{z}(z^{\prime}))$
Concept-based counterfactuals may also be explained via the interpretable model $f$. The following example illustrates a concrete scenario, also represented in Diagram 2.
Example. A CF-CBM address three questions by design: Predict concept $\hat{c}=\{\hat{c}_{yellow}=1,\hat{c}_{round}=0\}$ and class $\hat{y}=\{\hat{y}_{apple}=0,\hat{y}_{banana}=1\}$ labels. Explain that the class prediction is $\hat{y}_{banana}=1$ because $\hat{c}_{yellow}=1$ and $\hat{c}_{round}=0$ are relevant for the task predictor $f$. Imagine the counterfactual $\{\hat{c}_{yellow}^{\prime}=1,\hat{c}_{round}^{\prime}=1\}$ leading to the different class predictions $\hat{y}^{\prime}=\{\hat{y}_{apple}^{\prime}=1,\hat{y}_{banana}^{\prime}=0\}$.
3.4 Test-time Functionalities
In contrast to standard CBMs, CF-CBMs can either sample or estimate the most probable counterfactual to: (i) explain the effect of concept interventions on tasks, (ii) show users how to get a desired class label, and (iii) propose concept interventions via “task-driven” interventions.
Counterfactuals explain the effect of concept interventions on downstream tasks.
Plain concept-based explanations indicating the presence or absence of a concept for a given class prediction. However, the complexity of plain explanations grows quickly with the number of concepts. Concept-based counterfactuals, instead, induce simpler, sparser explanations (via Eq. 7) representing minimal modifications of concept labels that would have led to a different class prediction (Abid et al., 2022).
Act upon concept-based counterfactuals.
Concept-based counterfactuals can guide users towards achieving desired outcomes (indicated by class labels), especially when users do not know how to alter their status (represented by concepts). In this scenario: (i) a user specifies a desired class label $y^{\prime}$ for the downstream task representing their goal (e.g., get a loan), (ii) CF-CBMs generate a concept-based counterfactual $p(c^{\prime}|y^{\prime},\hat{y},\hat{c})$ representing minimal concept modifications changing the class label from the predicted $\hat{y}$ to the desired $y^{\prime}$ (e.g., save $10\%$ more each month), (iii) the user can act upon the counterfactual $c^{\prime}$ to accomplish the goal represented by $y^{\prime}$.
Task-driven interventions fix mispredicted concepts.
A key feature of CBMs is enabling human-in-the-loop interventions: at test time, users can correct mispredicted concept labels to enhance downstream task performance ($c\rightarrow y$ intervention). Concept interventions can be useful when intervening on concepts is easier than intervening on the downstream task. For instance, it might be easier to identify bone spurs ($c$) in a knee X-ray scan than to provide an accurate arthritis level ($y$). While still supporting concept interventions, CF-CBMs may also invert this mechanism via task-driven interventions ($y\rightarrow c$), exploiting its counterfactual generation abilities. Task-driven interventions can be used to correct mispredicted class labels when intervening on the downstream task is easier than intervening on concepts. For instance, it might be easier to identify a mispredicted Parkinson’s disease ($y$) rather than intervening in the genetic pathways ($c$) leading to the disease. In this scenario, users can intervene on the class label (rather than on the concepts) suggesting a correct $y^{\prime}$. Considering this additional information, CF-CBMs propose a more accurate set of concept labels $\hat{c}^{\prime}$ representing potential concept intervention previously unknown to the user.
Take the best bet or contemplate a multiverse?
All these functionalities can be used in two ways: the “best bet” mode and “multiverse” mode. In the “best bet” mode, CF-CBMs provide maximum a posteriori estimates by taking the mode of posterior distributions. Using this mode, users will be provided with the most probable counterfactual for each input. In “multiverse” mode instead, CF-CBMs sample from posterior distributions. This allows the model to generate multiple counterfactuals for each input, thus allowing users to choose among a variety of actions/interventions. For instance, in a healthcare scenario, the “best bet” mode might suggest the most probable alternative treatment plan based on patient data, while the “multiverse” mode could present a range of treatment strategies under varying clinical conditions, each representing a different plausible future.
4 Experiments
Our experiments aim to answer the following questions:
  • What? prediction generalization: Do CF-CBMs attain similar task/concept accuracy as standard CBMs?
  • Why? explanation quality: Can CF-CBMs produce simple and robust explanations?
  • What if? counterfactual actionability: Can CF-CBMs produce valid, plausible, efficient counterfactuals? Can CF-CBMs generate accurate task-driven interventions?
This section describes essential information about experiments. We provide further details in Appendix B.
4.1 Data & task setup
In our experiments we use three different datasets commonly used to evaluate CBMs: dSprites (Matthey et al., 2017), where the task is to predict specific combinations of objects having different shapes, positions, sizes and colours; MNIST addition (Manhaeve et al., 2018), where the task is to predict the sum of two digits; and CUB (Welinder et al., 2010), where the task is to predict bird species based on bird characteristics. In all experiments, images are encoded using a pre-trained ResNet18 (He et al., 2015).
4.2 Evaluation
In our analysis we use the following metrics. What? (prediction generalization): we compute the Area Under the Receiver Operating Characteristic Curve (Hand and Till, 2001) for concepts and tasks (ROC AUC ($\uparrow$)111($\uparrow$): the higher the better. ($\downarrow$): the lower the better.). Why? (explanation quality): we compute the impact of each concept on the downstream task, inspecting the task predictor $f$ (Yuksekgonul et al., 2022), and the impact of intervening on confounding concepts for class predictions (Koh et al., 2020). What if? (counterfactual actionability): drawing from previous works on counterfactuals, we compute: (i) the validity [^2] A.k.a. “correctness” (Abid et al., 2022).($\uparrow$(Wachter et al., 2017) by checking whether the model predicts the desired class labels based on the generated counterfactual; (ii) the proximity ($\downarrow$(Pawelczyk et al., 2020) which evaluates counterfactuals’ “reliability” as their similarity w.r.t. training samples; (iii) the time ($\downarrow$)  (Romashov et al., 2022) to generate a counterfactual for a set of samples.
In addition, we propose the following metrics to assess the quality of counterfactuals: (i) the $\Delta$-Sparsity ($\downarrow$) which evaluates user’s “wasted efforts” by counting the number concepts changed (Guo et al., 2023) w.r.t. the minimal number of changes that would have generated a counterfactual according to the dataset; (ii) the “plausibility” (Wachter et al., 2017) as the Intersection over Union (IoU ($\uparrow$)) (Jaccard, 1912) between counterfactuals and ground-truth concept vectors; (iii) the variability ($\uparrow$) which evaluates counterfactuals’ “diversity”, as the cardinality ratio between the set of counterfactuals generated and the set of training concept vectors; (iv) finally, we measure the concept accuracy of generated interventions (Acc Int. ($\uparrow$)) w.r.t. ground-truth optimal interventions that evaluate the model’s ability to generate potential concept interventions. Following Espinosa Zarlenga et al. (2022), we inject noise on predicted concepts to reduce task accuracy, then we sample counterfactuals, conditioning on the ground-truth label, and use them to fix mispredicted concept labels automatically.
All metrics are reported using the mean and the standard error over five different runs with different initializations.
4.3 Baselines
In our experiments, we compare CF-CBMs with standard CBMs in terms of generalization performance. To ensure a fair comparison, we design counterfactual baselines by adapting post-hoc methods to standard CBMs. We propose this experimental design for two main reasons. First, standard CBMs cannot generate counterfactuals by design, hence they need an external generator. Second, the standard, direct application of post-hoc methods on unstructured data types (e.g., image pixels rather than concepts) would have significantly compromised counterfactuals’ interpretability. We remark that this already represents an undocumented procedure in the current literature, as post-hoc methods typically derive counterfactuals from input features, not in the concept space. Among post-hoc approaches we select: (i) the Bayesian Counterfactual (BayCon) (Romashov et al., 2022), which generates counterfactuals via probabilistic feature sampling and Bayesian optimization; (ii) the Counterfactual Conditional Heterogeneous Variational AutoEncoder (C-CHVAE) (Pawelczyk et al., 2020), which iteratively perturbs the latent space of a VAE until it finds counterfactuals; (iii) the VAE CounterFactual (VAE-CF) (Mahajan et al., 2020), which improves C-CHVAE by conditioning counterfactuals on model’s predictions; (iv) the Variational Counter Net (VCNet) (Guyomard et al., 2023), which conditions latent counterfactual sampling on a specific target label $y^{\prime}$.
5 Key Findings & Results
5.1 What? Prediction Generalization
CF-CBMs match generalization performance of standard CBMs.
CF-CBMs attain both concept and task ROC AUC similar to standard CBMs (details about these results in Appendix C). This finding demonstrates that generating counterfactuals does not have a negative impact on classification performances on the datasets we considered. This outcome highlights how CF-CBMs generalize standard CBM architectures by matching their predictive accuracy while introducing new fundamental functionalities (i.e., counterfactual explanations and task-driven interventions).
5.2 Why? Explanation Quality
CF-CBMs provide simple and robust concept-based explanations for their predictions (Figure 3).
Following Koh et al. (2020), CF-CBMs provide concept-based explanations showing how each concept contributes to the downstream task prediction (Figure 3). To assess the impact of CF-CBMs design on the explanation’s quality we experiment using a modified version of the dSprites dataset, following Goyal et al. (2019). First, we introduced a high correlation between an object’s color and its class label. Then we removed this correlation on test images, and we checked whether the objects’ color was relevant for class predictions and whether intervening on objects’ colors impacted class predictions. We observed that standard CBM explanations considered both shapes and colors equally important for the class prediction, while CF-CBMs correctly focused on shapes only. Additionally, interventions on an object’s colors influenced $\sim 15\%$ of class predictions in standard CBMs, showing how the model was relying on confounders to solve the task via learning shortcuts (Marconato et al., 2023). CF-CBMs’ predictions, instead, were not affected by interventions on an object’s color, demonstrating better robustness to confounding factors. This robustness stems from CF-CBM’s design to generate a counterfactual altering the minimal number of concept labels to change the class label (e.g., changing the shape but not the color), leading the task predictor $f$ to prioritize the minimal set of decisive features. Appendix C provides further details.
(a) CBM
(b) CF-CBM
Figure 3: CF-CBM produce simpler and more robust concept-based explanations. The plots show two concept-based explanations representing the concept relevance given by the task predictor $f$ for the downstream task on the dSprites data set. Here, class labels depend on objects’ shapes only. Standard CBMs almost equally rely on both objects’ shapes and colors, while CF-CBMs mostly rely on shapes only.
5.3 What if? Counterfactual actionability
dSprites ($\uparrow$) MNIST add ($\uparrow$) CUB ($\uparrow$) avg. ($\uparrow$)
CBM+BayCon $\mathbf{100.0\pm 0.0}$ $94.6\pm 0.5$ $66.2\pm 0.5$ $86.9\pm 0.3$
CBM+CCHVAE $98.4\pm 0.6$ $94.7\pm 0.2$ $78.7\pm 0.4$ $90.6\pm 1.2$
CBM+VAECF $\mathbf{100.0\pm 0.0}$ $82.0\pm 5.9$ $95.2\pm 0.2$ $92.4\pm 2.0$
CBM+VCNET $\mathbf{100.0\pm 0.0}$ $89.9\pm 0.9$ $93.1\pm 0.5$ $94.3\pm 0.5$
CF-CBM (ours) $\mathbf{100.0\pm 0.0}$ $\mathbf{96.4\pm 0.7}$ $\mathbf{99.0\pm 0.1}$ $\mathbf{98.5\pm 0.3}$
Table 1: CF-CBMs outperform all baselines in counterfactuals’ validity across all datasets.
CF-CBMs outperform baselines in counterfactuals’ validity. (Table 1)
CF-CBMs attain the highest counterfactual validity across all datasets when compared to counterfactual generation baselines applied to the concept space of standard CBMs. This result shows that CF-CBMs can effectively generate concept vectors that lead to user-requested class labels. The gap w.r.t. baselines increases in complex datasets involving a vast search space over multiple concepts and classes. In the MNIST addition dataset (20 concepts, 19 classes), CF-CBMs achieve $+14$ percentage points (p.p.) higher validity than VAECF and $+7$ p.p. higher than VCNet. In CUB (118 concepts, 200 classes), CF-CBMs find solutions with a significantly higher validity w.r.t. iterative methods: $+33$ p.p. higher than BayCon and $+20$ p.p. higher than C-CHVAE.
Figure 4: CF-CBMs balance the trade off between counterfactuals’ reliability (proximity) and user effort ($\Delta$-Sparsity). The arrow ()normal-→($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region.
Figure 5: CF-CBMs calibrate the trade off between counterfactuals’ diversity (variability) and plausibility (IoU). The arrow ()normal-→($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region. “CF-CBM $n$” samples $n$ counterfactuals.
Figure 6: CF-CBMs generate reliable (proximity) and accurate interventions (ROC AUC Int.). The arrow ()normal-→($\rightarrow$)( → ) points towards optimal values. Pareto-optimal models form the frontier of the shaded region, whereas dominated solutions are located within the shaded region.
Figure 7: CF-CBMs efficiently generate counterfactuals on the fly, on par with the fastest generators.
CF-CBMs balance counterfactuals’ reliability with minimal user effort. (Figure 4)
High-quality counterfactuals should satisfy two key properties simultaneously: to be reliable and to require the least amount of users’ effort to act upon. However, minimizing user’s effort alone might lead to learning shortcuts producing less reliable counterfactuals, while optimizing reliability might generate counterfactuals that change the values of multiple concepts and represent a significant burden for users to take actions. For this reason, these two properties should be evaluated together:
  • CF-CBMs generate reliable counterfactuals: When processing large datasets, the training distribution covers a large portion of the feasible action space. For this reason, counterfactual solutions close to training samples are likely to be more reliable, as discussed by Pawelczyk et al. (2020); Laugel et al. (2019); Wachter et al. (2017). Our results show that CF-CBMs generate highly reliable counterfactuals across all datasets. Baseline models instead struggle to produce counterfactual close to training samples in at least one dataset.
  • CF-CBMs generate counterfactuals requiring minimal user effort to act upon: CF-CBMs consistently generate counterfactuals close to the optimal sparsity i.e., the fewest possible concept changes to alter the class prediction. This represents a significant result in terms of actionability because each concept modification requires user effort. For this reason, should a counterfactual require many concept label changes compared to the original concept vector, it would lead to a substantial unnecessary effort. Across all datasets, CF-CBMs stand out for their consistency in this aspect: their counterfactuals are either the sparsest or very close to the sparsest. This makes them highly effective and user-friendly, avoiding the burden of unnecessary actions.
CF-CBMs calibrate the variability-plausibility trade-off. (Figure 5)
Another key trade-off in generating counterfactuals corresponds to the tension between producing counterfactuals with high diversity, allowing users to pick among different roadmaps to action, and producing plausible counterfactuals corresponding to user actions that are likely to be feasible in practice. However, learning shortcuts might lead to trivial solutions with a poor variability-plausibility trade-off. For instance, once a plausible counterfactual is found, a trivial solution could be to generate always the same counterfacual, thus leading to minimal diversity, while maximizing the variability increases the chance of generating less plausible counterfactuals. In practice, these two properties are measured by variability and IoU between counterfactuals and ground-truth concept vectors. A higher variability implies that our model can produce a wider range of counterfactuals. A higher IoU value suggests that generated counterfactuals are also representative of the training distribution which contains plausible solutions. Our results show that CF-CBMs balance the optimization of these two properties better than existing methods. The gap w.r.t. BayCon and VCNet is already significant on dSprites where CF-CBMs produce counterfactuals with $+25$ p.p. higher IoU (plausibility) and ×$\times 3$ more variable. In MNIST addition the VAECF’s performances significantly deteriorate, while C-CHVAE mainly struggles in CUB.
CF-CBM (ours) CBM End-to-end NN BayCon C-CHVAE VAE-CF VCNet
What? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Why? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
What if? ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Concept interventions ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Task-driven interventions ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Real-time ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$ ${\color[rgb]{0,0.390625,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0.390625,0}\pgfsys@color@rgb@stroke{0}{0.390625}{0}\pgfsys@color@rgb@fill{0}{0.390625}{0}\bm{\checkmark}}$
Table 2: CF-CBMs efficiently address “what?”, “why?”, and “what if?” questions while supporting both concept and task-interventions, as opposed to existing CBMs and counterfactual generators.
CF-CBMs generate accurate task-driven interventions. (Figure 6)
CF-CBMs enable task-driven interventions showing users how to modify concepts to get the desired class label. This represents a novel crucial functionality when users are aware of a desired outcome but lack the knowledge required to intervene on concepts. CF-CBM significantly outperforms other baselines in concept accuracy of task-driven interventions by up to $78$ p.p. Such performance can be explained as the generation of task-driven interventions is conditioned on a user-provided target class label which provides fundamental information for the model to search in the latent concept distribution. Appendix C shows that CF-CBMs attain similar results while applying different levels of noise on predicted concept labels.
CF-CBMs efficiently generate counterfactuals on the fly. (Figure 7)
Generation efficiency is a key feature for real-time applications. In this respect, CF-CBMs significantly outperform iterative models such as C-CHVAE and BayCon. These models quickly struggle in large search spaces involving a high number of concepts: generating a counterfactual takes more than a minute on dSprites (7 concepts) and up to a few hours on CUB (118 concepts). CF-CBMs instead scale well with the size of the data set by requiring less than a second on dSprites and a few seconds on CUB, on par with VAECF and VCNet. This efficiency enables a wider range of applications requiring a high number of concepts and quick feedbacks.
6 Discussion & Conclusion
Limitations.
CF-CBMs combine features from CBMs and generative models, carrying forward their strengths and limitations. From generative models, CF-CBMs inherit the ability to efficiently approximate the data distribution. However, this approximation does not guarantee counterfactuals’ optimality when compared to exact methods (Wachter et al., 2017). Besides, CF-CBMs require concept annotations to ground explanations, as standard CBMs. Reasoning shortcuts (Marconato et al., 2023), concept impurities (Zarlenga et al., 2023), and information bottlenecks (Espinosa Zarlenga et al., 2022) are also typical limitations of CBM architectures, especially when the concept bottleneck is not complete (Yeh et al., 2020). However, recent works show how CBMs (and the same argument holds for CF-CBMs) could be extended using language models to approximate ground-truth concept annotations (Oikarinen et al., 2023) and concept embedding reasoners to provide robust task explanations reducing information bottlenecks (Barbiero et al., 2023).
Relations with concept-based models.
CF-CBMs share common ground with existing CBMs in modeling concepts in latent spaces. However, current models use the latent concept space to improve task accuracy (Espinosa Zarlenga et al., 2022; Barbiero et al., 2023), promote concept disentanglement (Misino et al., 2022; Marconato et al., 2022), or estimate confidence intervals for CBM’s predictions (Kim et al., 2023). Differently, CF-CBMs leverage the latent concept space to extend the capabilities of the CBM family by introducing two key functionalities (Table 2): counterfactual explanations and task-driven interventions.
Relations with generative structured models.
CF-CBMs are latent generative models characterized by a structured latent space. This characterization aligns them with other methodologies utilizing hierarchical priors (Sønderby et al. (2016); Maaløe et al. (2019); Vahdat and Kautz (2020); Klushyn et al. (2019)), mixture priors (Bauer and Mnih (2019)), and autoregressive priors (Chen et al. (2017); van den Oord et al. (2017); Razavi et al. (2019)). However, the structure of CF-CBMs’ latent variables explicitly captures both the concept and the counterfactual latent space, thus promoting structural interpretability, unlike existing structured latent variable models.
Relation with counterfactual models.
Existing DL methods generating counterfactuals typically offer post-hoc explanations (Goyal et al., 2019; Ghandeharioun et al., 2021; Abid et al., 2022) or require interpretable input features (Guyomard et al., 2023; Pawelczyk et al., 2020; Mahajan et al., 2020). However, applying post-hoc methods directly to unstructured data types (e.g., pixels rather than concepts) significantly reduces counterfactuals’ interpretability (Kim et al., 2018). In contrast, CF-CBMs: (i) are data-agnostic, allowing for the extraction of interpretable counterfactuals regardless of the data type, (ii) provide answers to “why?” questions by design via concept-based explanations, and (iii) support both concept and task-driven interventions (Table 2). Moreover, generating counterfactuals at the concept level brings significant advantages. First, concept-based counterfactuals are more interpretable and actionable, involving changes in high-level semantic units of information rather than raw features. As a side effect, concept-based counterfactuals may also introduce benefits in terms of privacy as sensitive and unique details embedded in input data (e.g., the image of a radiography) remain concealed, thus mitigating the risk of data extraction through repeated model queries (Goethals et al., 2023; Naretto et al., 2022; Pawelczyk et al., 2022).
Broader Impact
The ability to cover the ladder of interpretability represents a fundamental step to calibrate human trust and enhance human-machine interactions. In this work, we show how patching existing interpretable models (such as CBMs) and post-hoc counterfactual generators may represent a possible solution to address “what?”, “why?”, and “what if?” queries, thus covering the three rungs of the “ladder”. However, experimental results show that ad-hoc methods—designed and trained to address these three fundamental queries, such as CF-CBMs—may generate explanations and counterfactuals with higher quality. Besides, the ability to perform task-driven interventions may provide actionable insights to improve CF-CBMs’ predictions whenever users observe a mispredicted class label but do not know how to intervene on concepts. Finally, the proposed latent variable approach enables to efficiently model complex concept dependencies while generating multiple counterfactuals for a given input. This represents a key advantage w.r.t. iterative methods in practical applications where proposing a variety of interpretable counterfactuals on the fly represents a fundamental advantage, as in suggesting the quickest ways home to an autonomous vehicle or developing different treatment strategies under varying clinical scenarios.
7 Acknowledgments
This study was funded by TRUST-ME (project 205121L_214991), SmartCHANGE (GA No. 101080965) and XAI-PAC (PZ00P2_216405) projects.
Appendix A Ammortized inference and ELBO Derivation
To derive the ELBO objective function defined in Section 3.2, we start from the maximization of the log-likelihood of the tuple $(c,y,y^{\prime})$:
[…]
where $p(c,c^{\prime},y,y^{\prime},z,z^{\prime})$ factorizes as in Section 3.2. We consider the variational approximation $q(z,z^{\prime},c^{\prime}|x,y,c,y^{\prime})=q(z|x)q(z^{\prime}|z,c,y,y^{\prime})p(c^{\prime}|z^{\prime})$ thus exploiting the observation of $x$ to better infer $z$:
$x$$c$$y$$y^{\prime}$$z$$z^{\prime}$
We have:
$\log p(c,y,y^{\prime})=\log\int$ $\frac{q(z|x)q(z^{\prime}|z,c,y,y^{\prime})p(c^{\prime}|z^{\prime})}{q(z|x)q(z^{\prime}|z,c,y,y^{\prime})p(c^{\prime}|z^{\prime})}\cdot$
$\quad\quad\cdot p(c,c^{\prime},y,y^{\prime},z,z^{\prime})dc^{\prime}dzdz^{\prime}$
and, given the Jensen’s inequality, we obtain:
$\log p(c,y,y^{\prime})\geq\int$ $q(z|x)q(z^{\prime}|z,c,y,y^{\prime})p(c^{\prime}|z^{\prime})\cdot$
$\cdot\log\frac{p(c,c^{\prime},y,y^{\prime},z,z^{\prime})}{q(z|x)q(z^{\prime}|z,c,y,y^{\prime})p(c^{\prime}|z^{\prime})}dc^{\prime}dzdz^{\prime}$
that can be rewritten as:
$\log p(c,y,y^{\prime})\geq$ $\mathbb{E}_{z}[\log(c|z)]+$
$\log p(y|c)+$
$\mathbb{E}_{z,z^{\prime},c^{\prime}}[\log p(y^{\prime}|c^{\prime})]-$
$D_{KL}[p(z)|q(z|x)]-$
$D_{KL}[p(z^{\prime})|q(z^{\prime}|z,c,y,y^{\prime})].$
Appendix B Experimental details
B.1 Data & task setup
In our experiments we use three different datasets commonly used to evaluate CBMs:
  • dSprites [Matthey et al., 2017] — It is composed of images of one of three objects (square, ellipse, heart) in different positions and of different sizes. Starting from it, we design a binary task, which is to predict if in the image there is at least a square or a heart. We also combine initial images to obtain new ones with more than one object. The concepts for this task are: (1) the presence of a square, (2) the presence of an ellipse, (3) the presence of a heart, (4) the presence of two objects, (5) objects are red, (6) objects are green, (7) objects are blu.
  • MNIST addition — The task of this dataset is to predict the sum of two MNIST digits [Deng, 2012]. The concepts of this task are the one-hot encoding of the first and the second digits concatenated together.
  • CUB [Welinder et al., 2010] — It contains pictures of birds and the final task is to predict their species (among 200 species). The concepts are 118 bird features that human annotators selected.
B.2 Evaluation Metrics
In our analysis we use the following metrics. What? (prediction generalization): we compute the Area Under the Receiver Operating Characteristic Curve [Hand and Till, 2001] for concepts and tasks (ROC AUC ($\uparrow$). Why? (explanation quality): we compute the impact of each concept on the downstream task, inspecting the task predictor $f$ [Yuksekgonul et al., 2022]. For each class, we look at the related weights of the task predictor (a Linear layer), one weight for each concept. In this way, it is visible the impact of each concept on each specific class. Then, we evaluate the impact of intervening on confounding concepts for class predictions [Koh et al., 2020]. We modify the dSprites task to highly correlate the colors with the shapes and therefore the task label. For instance, in 85% of samples with a positive label (presence of at least one square or heart), the objects are green, while in 85% of the samples with a negative label, the objects are red or blue. What if? (counterfactual actionability): drawing from previous works on counterfactuals, we compute: (i) the validity($\uparrow$[Wachter et al., 2017] by checking whether the model predicts the desired class labels based on the generated counterfactual; (ii) the proximity ($\downarrow$[Pawelczyk et al., 2020] which evaluates counterfactuals’ “reliability” as their similarity w.r.t. training samples; (iii) the time ($\downarrow$)  [Romashov et al., 2022] to generate a counterfactual for a set of samples.
In addition, we propose the following metrics to assess the quality of counterfactuals: (i) the $\Delta$-Sparsity ($\downarrow$) which evaluates user’s “wasted efforts” by counting the number concepts changed [Guo et al., 2023] w.r.t. the minimal number of changes that would have generated a counterfactual according to the dataset:
$\Delta Sparsity=|OptimalSparsity-Sparsity|\$
where $OptimalSparsity$ represents the mean Hamming distance between each concept vector in the test set and the closest concept vector of a drawn random class $y^{\prime}$, found using a “brute-force” approach and $Sparsity$ represents the mean Hamming distance between the predicted concepts $\hat{c}$ and the predicted concept counterfactuals $\hat{c^{\prime}}$; (ii) the “plausibility” [Wachter et al., 2017] as the Intersection over Union (IoU ($\uparrow$)) [Jaccard, 1912] between counterfactuals and ground-truth concept vectors; (iii) the variability ($\uparrow$) which evaluates counterfactuals’ “diversity”, as the cardinality ratio between the set of counterfactuals generated and the set of training concept vectors
$Variability=\frac{|c^{\prime}|}{|c|}$
where $|c^{\prime}|$ represents the cardinality of the set composed by all $\hat{c^{\prime}}$, while $|c|$ represents the cardinality of the set composed by all $c$; (iv) finally, we measure the concept accuracy of generated interventions (Acc Int. ($\uparrow$)) w.r.t. ground-truth optimal interventions that evaluate the model’s ability to generate potential concept interventions. Following Espinosa Zarlenga et al. [2022], we inject noise on predicted concepts $\hat{c}$ to reduce task accuracy (randomly flip the value of some concepts), then we sample counterfactuals, conditioning on the ground-truth label $y$, and use them to fix mispredicted concept labels automatically.
$AccInt.=\frac{\sum_{i\in c}1_{c_{i}=\hat{c^{\prime}_{i}}}}{|c|}$
where $c_{i}$ represents the concept ground truth for the sample $i$, while $\hat{c^{\prime}_{i}}$ represents the counterfactual generated for the sample $i$ that tries to recover $c$.
B.3 Baselines and Implementation details
Hyperparameters
To train our models and the baselines we selected the best hyperparameters according to our experiments. Table 3 shows the number of epochs, learning rate, and embedding size in the latent space, batch size for each dataset. They are shared among the baselines, and we took the best checkpoint out of the entire training for each model. In addition, Table 4 illustrates the parameters used to weight each term in the loss. However, all these details are in the released code.
Hyperparams dSprites MNIST add CUB
Epochs 75 150 150
Learning rate 0.005 0.005 0.005
Hidden size 128 128 128
Batch size 1024 1024 1024
Table 3: Model Hyperparameters shared across all models. During the training, we select the best checkpoint for each model.
Method Dataset
Loss Term dSprites MNIST add CUB
CBM Concept 1.0 1.0 1.0
Task 0.1 0.1 0.1
CCHVAE Reconstruction 1.0 1.0 1.0
KL 0.5 0.5 0.5
VAECF Reconstruction 3.0 1.3 1.0
KL 1.0 1.0 2.0
Validity 15.0 15.0 15.0
VCNET Concept 1.0 1.0 1.0
Task 0.5 0.5 0.5
Reconstruction 0.5 0.5 0.5
KL 0.8 0.8 0.8
CF-CBM Concept 10.0 10.0 1.0
Task 0.7 1.0 0.1
Validity 0.3 0.2 0.02
$z$ KL 1.2 2.0 0.2
$z^{\prime}$ KL 1.2 2.0 0.2
Prior distance 1.0 1.7 0.2
Posterior distance 0.6 0.55 0.03
Table 4: Loss weights
Counterfactual CBM implementation details
To improve the training process of our models we decided to relax some assumptions we made in Section 3.2. Employing a Bernoulli distribution to predict concepts and a Categorical distribution for the class prediction would make the training process more difficult, ruining the gradient in the backpropagation process. Therefore, we predict them directly with the output of the concept predictor and the task predictor applying a Sigmoid function and a Softamx on top of them, respectively. To optimise them, we choose to use a Binary Cross Entropy loss for the concept loss, and a Cross Entropy Loss for the task loss.
Code and licenses
For our experiments, we implement all baselines and methods in Python 3.9 and relied upon open-source libraries such as PyTorch 2.0 [Paszke et al., 2019] (BSD license), and Sklearn 1.2 [Pedregosa et al., 2011] (BSD license). In addition, we used Matplotlib [Hunter, 2007] 3.7 (BSD license) to produce the plots shown in this paper. We will publicly release the code with all the details used to reproduce all the experiments under an MIT license.
Appendix C Further metrics and results
In addition to the result present in Section 5, Table 5 shows the task and concept performance of our model and the baselines. Moreover, Figure 8 illustrates the Acc Int. achieved by all models with different levels of noise injected at the concept level. It is visible how CF-CBM achieves stable Acc Int. results with different level of noise injected at the concept level. This is not the case for the other baselines. Finally, in contrast to Section 5, which shows an example of global feature importance for the dSprites dataset, Figure 9 represents the local feature importance of our model. Specifically, it illustrates how the model assesses the importance of features for predicting a specific class based on a single sample from the dataset.
dsprites MNIST add CUB
Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$) Task (%) ($\uparrow$) Concept (%) ($\uparrow$)
CBM $99.6\pm 0.1$ $98.8\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.0\pm 0.0$ $85.3\pm 0.1$
BayCon $99.6\pm 0.0$ $98.7\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.1\pm 0.1$ $85.3\pm 0.1$
CCHVAE $99.5\pm 0.1$ $98.9\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $93.2\pm 0.1$ $85.4\pm 0.1$
VAECF $99.5\pm 0.0$ $98.8\pm 0.1$ $98.4\pm 0.0$ $99.7\pm 0.0$ $92.9\pm 0.2$ $85.4\pm 0.1$
VCNET $99.7\pm 0.0$ $98.6\pm 0.1$ $98.2\pm 0.0$ $99.6\pm 0.0$ $89.6\pm 0.5$ $80.7\pm 0.1$
ACBM $99.7\pm 0.0$ $98.7\pm 0.1$ $97.0\pm 0.1$ $99.3\pm 0.0$ $91.4\pm 0.2$ $84.8\pm 0.2$
Table 5: CF-CBM attains generalization performance of standard CBMs on concepts and class labels.
Figure 8: CF-CBMs attain similar results while applying different levels of noise on predicted concept labels.
Figure 9: The plots show the concept-based explanations representing the concept relevance given by the task predictor $f$ for a specific class of the downstream task data set. The importance of each concept for the task prediction is highlighted using color-coded boxes: red for positive importance and blue for negative impact. The size of each box correlates with the magnitude of the concept’s impact. The numerical values at the bottom of each box denote the activation value of the respective concept.